# Dim and Small Target Detection for Drone Broadcast Frames Based on Time-Frequency Analysis

### v1

@@ Abstract
We propose a dim and small target detection algorithm for drone broadcast frames based on the time-frequency analysis of communication protocol. Specifically, by analyzing modulation parameters and frame structures, the prior knowledge of transmission frequency, signal bandwidth, Zadoff-Chu (ZC) sequences, and frame length of drone broadcast frames is established. The RF signals are processed through the designed filter banks, and the frequency domain parameters of bounding boxes generated by the detector are corrected with transmission frequency and signal bandwidth. Given the remarkable correlation characteristics of ZC sequences, the frequency domain parameters of bounding boxes with low confidence scores are corrected based on ZC sequences and frame length, which improves the detection accuracy of dim targets under low signal-to noise ratio (SNR) situations. Besides, a segmented energy refinement method is applied to mitigate the deviation caused by interference signals with high energy strength, which ulteriorly corrects the time domain detection parameters for dim targets. As the sampling duration increases, the detection speed improves while the detection accuracy of broadcast frames termed as small targets decreases. The trade-off between detection accuracy and speed versus sampling duration is established, which helps to meet different drone regulation requirements. Simulation results demonstrate that the proposed algorithm improves the average intersection over union, precision, and recall by 3%, 1.4%, and 2.4%, respectively, compared to existing algorithms. The proposed algorithm also performs strong robustness under varying flight distances, diverse types of environment noise, and different flight visual environment.
@@ I Introduction
The low-altitude economy, which is rapidly developing, refers to an economic model focused on the flight activities of both manned and unmanned aerial vehicles within airspace ranging from the ground up to 1000 m[1]. The low-altitude system composed of low-cost and highly flexible drones, serves as the foundational infrastructure for the low-altitude industrialization and a key component of the low-altitude economy[2]. While the widely deployed drones have enhanced economic growth across various areas including logistics, agriculture, and tourism, they have also introduced new security threats, such as unauthorized malicious flights and eavedropping attacks[3],[4],[5],[6]. Although many countries and organizations have issued policies and rules for drone registration, operator training, and flight authorization, there remains a lack of effective methods for drones remote identification (RID)[7]. Existing works mainly focus on the detection and classification of drones, which have neglected the object tracking and trajectory prediction. Specifically, the radar echo signals[8],[9], visual images[10], acoustic signals[11], and radio-frequency (RF) signals[12],[13] are commonly utilized to detect and classify the types and operation modes of drones. However, effective drone regulation requires accurate tracking and prediction of drone positions and trajectories, as mere detection and classification are insufficient for further analyzing drone intents and implementing countermeasures.
@@ I-B Motivations and Contributions
  • Research on utilizing drone broadcast frames to achieve RID remains insufficient. Unlike non-communication protocol-based algorithms, broadcast frames transmit plain text information as required by communication protocols, which provides a promising choice for drones tracking and prediction. Thus, the drone RID problem can be transformed into two key issues: drone detection and classification in[16], and the detection and decoding of broadcast frames.
@@
  • For the detection of broadcast frames termed as dim and small targets, data-level decoupling is more effective and accurate than model-level decoupling. Existing image reconstruction and enhancement algorithms for dim and small target object detection, such as dehazing and denoising, fail to consistently and reliably improve detection accuracy of signal broadcast frames. This is because interference signals and broadcast frames exhibit similar characteristics in TFIs, and pixel-level image reconstruction may not aid semantic-level target detection, which even degrades the accuracy of FHSS signals and broadcast frames under low SNR and signal overlapping situations. Nevertheless, by analyzing communication protocols to obtain prior knowledge of modulation parameters and frame structures, data-level decoupling can be adopted to improve the model’s detection accuracy when detecting broadcast frames termed as dim and small targets.
@@
  • In practical environments, detecting broadcast frames termed as dim and small targets in TFIs often leads to bounding boxes significantly deviating from the ground truth, which means the discrepancies in time-frequency parameters. Under low SNR situations, broadcast frames may be masked by background noise, causing bounding boxes to appear at non-target time-frequency positions with low confidence scores. When interference signals exhibit high energy strength or overlap with broadcast frames, bounding boxes often appear in the time-frequency regions of interference signals with low confidence scores. Since bounding boxes with high confidence scores tend to have relatively accurate detection positions, it is necessary to correct the time and frequency domain parameters of bounding boxes with low confidence scores, which will improve the detection accuracy of broadcast frames termed as dim and small targets.
  • The trade-off between the detection speed and accuracy of broadcast frames needs to be carefully considered. Specifically, unlike other types of object images, TFIs with a fixed size of $H\times W\times 3$ must maintain a bandwidth of approximately 100 MHz due to communication protocol constraints. However, samples with different sampling durations reveal varying levels of detail information, which means broadcast frames do not always appear as dim and small targets in TFIs. While shorter sampling durations can improve detection accuracy, the detection speed of pre-trained detector, such as frame per second (FPS) or detection latency, remains fixed. Therefore, it is essential to carefully balance the trade-off between FPS[45] or latency requirements[11],[12],[16],[26] and detection accuracy.
@@
To address the above challenges, we propose a time-frequency analysis-based dim and small target detection algorithm for drone broadcast frames, utilizing outdoor practical sampling datasets named DroneRFa[46] and DroneRFb-DIR[47]. Specifically, the YOLOv7 network[48], a widely used object detection model, is leveraged to identify and locate broadcast frames within TFIs, while the prior knowledge of modulation parameters and frame structures is derived from communication protocols. The frequency domain parameters of bounding boxes are corrected by the transmission frequency and signal bandwidth, while the time-domain parameters are modified based on the cross-correlation results of Zadoff-Chu (ZC) sequences and frame length. Additionally, the impact of interference signals with high energy strength on cross-correlation results is mitigated by adopting a segmented energy refinement method. The contributions of this paper are summarized as follows.
@@
  • This paper acquires prior knowledge of the modulation parameters and frame structures of drone broadcast frames, and the data-level decoupling is performed within the detection architecture. The potential transmission frequencies of broadcast frames, spanning the 100 MHz sampling bandwidth at 2.4 GHz and 5.8 GHz as regulated by communication protocols, are statistically summarized. Additionally, based on the estimation of signal bandwidth and the number of subcarriers, the ZC sequences adopted in broadcast frames are identified through an analysis of the signal frame structure.
@@
  • Leveraging prior knowledge, a data-level decoupling algorithm is proposed to correct the time and frequency domain parameters of bounding boxes. Three sets of filter banks, specifically designed for 2.4 GHz and 5.8 GHz based on the transmission frequency set and signal bandwidth, are applied before detection. For bounding boxes within the filtered TFIs, frequency domain parameters are corrected according to the fixed signal bandwidth of broadcast frames. Motivated by the remarkable correlation properties of ZC sequences, the ZC sequences adopted in broadcast frames are locally generated to enable accurate detection under low SNR and interference situations. The peak’s index of the ZC sequence cross-correlation results is utilized to determine the beginning location of broadcast frames, while the time domain parameters of low-confidence bounding boxes are corrected based on the prior knowledge of frame length. Furthermore, to mitigate the impact of interference signals with high energy strength on cross-correlation results, a segmented energy refinement module is deployed.
@@
  • The trade-off between detection accuracy and speed of the proposed drone broadcast frame detection algorithm is established. Although the FPS of the YOLOv7 detector is fixed, detection latency varies with different sampling durations of RF signals due to the cost of time domain parameters correction. Additionally, detection accuracy fluctuates since varying sampling durations affect whether broadcast frames appear as dim and small targets in TFIs. By performing broadcast frame detection for RF signals with different sampling durations, the trade-off between detection accuracy and speed of the proposed algorithm is established, which helps to meet different detection requirements of drone regulation.
@@
  • Numerical results demonstrate that the proposed algorithm effectively achieves the dim and small target detection for drone broadcast frames. Specifically: 1) The correction of bounding boxes based on time-frequency analysis of communication protocol is highly effective, and data-level decoupling enhances detection accuracy compared to model-level decoupling in the presence of interference signals and background noise. The proposed algorithm achieves improvements of 3%, 1.6%, and 2.4% in intersection over union (IoU), precision, and recall, respectively. 2) The proposed algorithm exhibits strong robustness across varying flight distances, diverse types of noise, different flight visual environment, which maintains approximately 70% detection accuracy under the poor channel conditions. 3) The proposed algorithm can achieve the best trade-off between detection accuracy and speed under different drone detection scenarios, with a maximum detection accuracy, latency, and speed of 90%, 100 ms, and 55 FPS, respectively.
@@ I-C Outline and Notations
The signal analysis is presented in Section II. The details and explanations of drone broadcast frames detection algorithm are discussed in Section III. Simulation evaluation is presented in Section IV. And this paper is summarized in Section V.
@@ II-C Estimation of Modulation Parameters
    Specifically, for the estimated signal bandwidth $\hat{B}_{u}=9.015$ MHz, given the corresponding subcarrier interval of 15 KHz, it can be calculated that there exists $N_{u}=600$ used subcarriers and 1 direct-current subcarrier. Then, it can also be inferred that there are $N_{v}-1$ virtual subcarriers on both sides of the used subcarriers, with 212 on one side and 211 on the other side. For OFDM signals with different $N$ and $B$, the bandwidth estimation is performed after sampling with $f_{s}$, and the simulation results are shown in Fig. 3. The parts per million (PPM) can be computed by (6), and it can be observed that the estimation error of $B$ does not exceed 120 KHz for commonly used $N=1024$ and 2048. Moreover, since OFDM signals have a fixed set of $B$ in MHz, such an error has no impact on acquiring prior knowledge of the signal bandwidth of broadcast frames.
    $\text{PPM}=\frac{\left|B-\hat{B}\right|}{10^{6}}.$ (6)
@@
    Number of Subcarriers: Since OFDM signals add cyclic prefix (CP) before each symbol to eliminate inter-symbol interference and inter-carrier interference caused by multipath propagation, the auto-correlation characteristic of CP can be utilized to estimate the number of subcarriers $N$. The auto-correlation results for $x(l)$ with a window length $L_{a}$ can be obtained by
@@ III-A Frequency Domain Parameters Correction
$\left\{\begin{array}[]{cc}\hat{y}_{\min}=\frac{y_{\max}+y_{\min}}{2}-\frac{B}{2},&\frac{y_{\max}+y_{\min}}{2}>f_{i}\\ \hat{y}_{\min}=f_{i}-\frac{B}{2},&\frac{y_{\max}+y_{\min}}{2}\leq f_{i},\end{array}\right.$ (14)
@@
Figure 7: Bouding box with confidence score of T0101D00S00L1 at SNR = -7 dB. (a) Without frequency domain parameters correction. (b) With frequency domain parameters correction.
Figure 8: Bouding box with confidence score of T0011D00S00L4 at SNR = 15 dB. (a) Original Bounding Box. (b) With frequency domain parameters correction. (c) With time domain parameters correction.
@@ III-B Time Domain Parameters Correction
Due to the presence of noise and interference signals, the time domain parameters of the bounding boxes may deviate from the ground truth. The detected duration of the drone broadcast frame can be expressed as
$T_{d}=x_{\max}-x_{\min},$ (15)
while the duration of drone broadcast frames regulated can be computed by
$T=\frac{M(N+N_{\text{cp}})}{B},$ (16)
where $M$ denotes the number of OFDM symbols in one broadcast frame. It is worth noting that, according to the prior knowledge of the communication protocol, $M$ and $N_{\text{cp}}$ will vary based on the drone type and the OFDM symbol index.
@@
For time domain parameter deviation caused by the presence of interference signals, as shown in Fig. 9a, the correction of $x_{\min}$ and $x_{\max}$ can be performed directly based on (17) and (18). However, under low SNR situations, bounding boxes may be generated at the time-frequency positions of interference signals with high energy strength, as shown in Fig. 10a. Although $y_{\min}$ and $y_{\max}$ can be corrected by (13) and (14), the correction of $x_{\min}$ and $x_{\max}$ based on (17) and (18) merely will result in the wrong $m^{*}$ corresponding to the interference signals, as shown in Fig. 11a. Thus, a segmented energy refinement method is applied to mitigate the time domain parameters deviation caused by the interference signals with high energy strength. $\gamma(m)$ will be divided into $Q$ non-overlapping sequences with length $P$, and the mean energy of each segment can be calculated by
@@
Figure 9: Bouding box with confidence score of T0011D00S00L0 at SNR = 15 dB. (a) Without time domain parameters correction. (b) With time domain parameters correction.
Figure 10: Bouding box with confidence score of T0001D00S00L1 at SNR = -9 dB. (a) Original Bounding Box. (b) With frequency domain parameters correction. (c) With time domain parameters correction.
@@
IV Numerical Results
IV-A Simulation Setup

### v2

@@ Abstract
We propose a dim and small target detection algorithm for drone broadcast frames based on the time-frequency analysis of communication protocol. Specifically, by analyzing modulation parameters and frame structures, the prior knowledge of transmission frequency, signal bandwidth, Zadoff-Chu (ZC) sequences, and frame length of drone broadcast frames is established. The RF signals are processed through the designed filter banks, and the frequency domain parameters of bounding boxes generated by the detector are corrected with transmission frequency and signal bandwidth. Given the remarkable correlation characteristics of ZC sequences, the frequency domain parameters of bounding boxes with low confidence scores are corrected based on ZC sequences and frame length, which improves the detection accuracy of dim targets under low signal-to noise ratio situations. Besides, a segmented energy refinement method is applied to mitigate the deviation caused by interference signals with high energy strength, which ulteriorly corrects the time domain detection parameters for dim targets. As the sampling duration increases, the detection speed improves while the detection accuracy of broadcast frames termed as small targets decreases. The trade-off between detection accuracy and speed versus sampling duration is established, which helps to meet different drone regulation requirements. Simulation results demonstrate that the proposed algorithm improves the evaluation metrics by 2.27% compared to existing algorithms. The proposed algorithm also performs strong robustness under varying flight distances, diverse types of environment noise, and different flight visual environment. Besides, the broadcast frame decoding results indicate that 97.30% accuracy of RID has been achieved.
@@ I Introduction
The low-altitude economy, which is rapidly developing, refers to an economic model focused on the flight activities of both manned and unmanned aerial vehicles within airspace ranging from the ground up to 1000 m[1]. The low-altitude system composed of low-cost and highly flexible drones, serves as the foundational infrastructure for the low-altitude industrialization and a key component of the low-altitude economy[2]. While the widely deployed drones have enhanced economic growth across various areas including logistics, agriculture, and tourism, they have also introduced new security threats, such as unauthorized malicious flights and eavedropping attacks[3],[4],[5],[6]. Although many countries and organizations have issued policies and rules for drone registration, operator training, and flight authorization, there remains a lack of effective methods for drones remote identification (RID)[7]. Existing works mainly focus on the detection and classification of drones, which have neglected the object tracking and trajectory prediction. Specifically, the radar echo signals[8],[9], visual images[10], acoustic signals[11], and radio-frequency (RF) signals[12],[13] are commonly utilized to detect and classify the types and operation modes of drones. However, effective drone regulation requires accurate tracking and prediction of drone positions and trajectories, as mere classification are insufficient for further analyzing drone intents and implementing countermeasures.
@@ I-B Motivations and Contributions
  • Insufficient research on leveraging drone broadcast frames for RID. Unlike non-protocol-based algorithms, broadcast frames convey plain text information as required by communication protocols, providing a promising choice for drone tracking and trajectory prediction. Thus, the tracking and prediction task can be transformed into the detection and decoding of broadcast frames.
@@
  • Superiority of data-level over model-level decoupling for broadcast frame detection termed as dim and small targets. Traditional image enhancement methods, such as dehazing or denoising, used in dim and small target detection are ineffective for broadcast frames. This is due to the similar characteristics between broadcast frames and interference signals in TFIs. Pixel-level image enhancements may not consistently and reliably improve semantic-level detection, especially under low-SNR and signal overlapping conditions. In contrast, leveraging prior knowledge of communication protocol, such as modulation parameters and frame structures, enables data-level decoupling and detection performance enhancement.
@@
  • Bounding box deviations in TFIs under practical environments. Under low SNR situations, broadcast frames may be obscured by noise, causing bounding boxes to deviate significantly from the ground truth. In the presence of high-energy interference or overlapping signals, bounding boxes may shift to the time-frequency regions of the interference, often with low confidence scores. Since bounding boxes with high confidences are typically more accurate, it becomes necessary to correct the time-frequency parameters of those with low confidence to improve overall detection precision.
  • Trade-off between detection accuracy and speed in broadcast frame detection. Unlike natural image-based object detection tasks, TFIs used for drone RID must retain a bandwidth of approximately 100 MHz to conform to protocol constraints. While different sampling durations influence whether broadcast frames appear as dim and small targets in TFIs, the inference speed (e.g., frame per second (FPS)[45] or latency[11],[12],[16],[26]) of the pre-trained detector remains fixed. Thus, optimizing the trade-off between sampling duration and detection accuracy is crucial for meeting real-time RID requirements.
@@
To address the above challenges, we propose a time-frequency analysis-based dim and small target detection algorithm for drone broadcast frames, utilizing outdoor practical sampling datasets named DroneRFa[46] and DroneRFb-DIR[47]. Specifically, the YOLOv7 network[48] is leveraged to identify and locate broadcast frames within TFIs, while the prior knowledge of modulation parameters and frame structures is derived from communication protocols. The frequency domain parameters of bounding boxes are corrected by the transmission frequency and signal bandwidth, while the time-domain parameters are modified based on the cross-correlation results of Zadoff-Chu (ZC) sequences and frame length. Additionally, the impact of interference signals with high energy strength on cross-correlation results is mitigated by adopting a segmented energy refinement method. The contributions of this paper are summarized as follows.
@@
  • Prior knowledge of the modulation parameters and frame structures of drone broadcast frames is systematically established. The potential transmission frequencies of broadcast frames, spanning the 100 MHz sampling bandwidth at 2.4 GHz and 5.8 GHz as regulated by communication protocols, are statistically summarized. Additionally, the ZC sequences adopted in broadcast frames are identified based on the estimation of signal bandwidth and the number of subcarriers.
@@
  • A data-level decoupling algorithm is proposed for bounding box parameter correction based on protocol prior knowledge. Three sets of filter banks, specifically designed for 2.4 GHz and 5.8 GHz based on the transmission frequency set and signal bandwidth, are applied before detection. For bounding boxes in the filtered TFIs, frequency domain parameters are corrected according to the fixed signal bandwidth. The peak’s index of the ZC sequence cross-correlation results is utilized to determine the beginning location of broadcast frames, while the time domain parameters of low-confidence bounding boxes are corrected based on the prior knowledge of frame length. Furthermore, to mitigate the impact of interference signals with high energy strength on cross-correlation results, a segmented energy refinement module is deployed.
@@
  • The trade-off between detection accuracy and speed is quantitatively analyzed. Although the YOLOv7 detector’s FPS remains fixed, detection latency varies depending on the sampling duration due to time domain parameters correction overhead. Detection accuracy fluctuates since sampling duration affects whether broadcast frames appear as dim and small targets in TFIs. By evaluating broadcast frame detection performance across different sampling durations, the proposed algorithm provides flexibility for adapting to various drone regulation scenarios.
  • A decoding algorithm for broadcast frames has been proposed. Based on the time-frequency parameters of the bounding boxes, cyclic prefix (CP) and ZC sequences are utilized for time and frequency synchronization, and bitstream is obtained. Based on the analysis of communication protocols, the descrambling and Turbo decoding are performed, and the payload information is recovered according to the regulated message format.
@@
  • Numerical results demonstrate that the proposed algorithms effectively achieve drone RID across different channel conditions and drone types. Specifically: 1) The proposed algorithm achieves at least 3%, 1.6%, and 2.4% improvement in intersection over union (IoU), precision, and recall over existing baselines, confirming the efficacy of time-frequency analysis of communication protocols. 2) The algorithms exhibit strong robustness across variable flight distances, visual environments, and noise types. 3) The proposed algorithms achieve the best trade-off between detection accuracy and latency, ensuring feasibility for practical deployment. 4) The decoding results for different drone broadcast frames indicate that the proposed algorithm achieve effective RID.
@@ I-C Outline and Notations
The signal analysis is presented in Section II. The details of drone broadcast frames detection and decoding algorithm are discussed in Section III and Section IV, respectively. Simulation evaluation is presented in Section V. And this paper is summarized in Section VI.
@@ II-C Estimation of Modulation Parameters
    Specifically, for the estimated signal bandwidth $\hat{B}_{u}=9.015$ MHz, given the corresponding subcarrier interval of 15 KHz, it can be calculated that there exists $N_{u}=600$ used subcarriers and 1 direct-current subcarrier. Then, it can also be inferred that there are $N_{v}-1$ virtual subcarriers on both sides of the used subcarriers, with 212 on one side and 211 on the other side. For OFDM signals with different $N$ and $B$, the bandwidth estimation is performed after sampling with $f_{s}$, and the simulation results are shown in Fig. 3. The parts per million (PPM) can be computed by $\frac{\left|B-\hat{B}\right|}{10^{6}}$, and it can be observed that the estimation error of $B$ does not exceed 120 KHz for commonly used $N=1024$ and 2048. Moreover, since OFDM signals have a fixed set of $B$ in MHz, such an error has no impact on acquiring prior knowledge of the signal bandwidth of broadcast frames.
@@
    Number of Subcarriers: Since OFDM signals add CP before each symbol to eliminate inter-symbol interference and inter-carrier interference caused by multipath propagation, the auto-correlation characteristic of CP can be utilized to estimate the number of subcarriers $N$. The auto-correlation results for $x(l)$ with a window length $L_{a}$ can be obtained by
@@ III-A Frequency Domain Parameters Correction
Figure 7: Bouding box of T0101D00S00L1 at SNR = -7 dB. (a) Original Bounding Box. (b) With frequency domain parameters correction.
Figure 8: Bouding box of T0011D00S00L4 at SNR = 15 dB. (a) Original Bounding Box. (b) With frequency domain correction. (c) With time domain correction.
@@ III-B Time Domain Parameters Correction
Due to the presence of noise and interference signals, the time domain parameters of the bounding boxes may deviate from the ground truth. The detected duration of the drone broadcast frame is $T_{d}=x_{\max}-x_{\min}$, while the regulated duration can be computed by $T=\frac{M(N+N_{\text{cp}})}{B}$, where $M$ denotes the number of OFDM symbols in one broadcast frame. It is worth noting that, according to the prior knowledge of the communication protocol, $M$ and $N_{\text{cp}}$ will vary based on the drone type and the OFDM symbol index.
@@
For time domain parameter deviation caused by the presence of interference signals, as shown in Fig. 9a, the correction of $x_{\min}$ and $x_{\max}$ can be performed directly based on (13). However, under low SNR situations, bounding boxes may be generated at the time-frequency positions of interference signals with high energy strength, as shown in Fig. 10a. Although $y_{\min}$ and $y_{\max}$ can be corrected by (11), the correction of $x_{\min}$ and $x_{\max}$ based on (13) merely will result in the wrong $m^{*}$ corresponding to the interference signals, as shown in Fig. 11a. Thus, a segmented energy refinement method is applied to mitigate the time domain parameters deviation caused by the interference signals with high energy strength. $\gamma(m)$ will be divided into $Q$ non-overlapping sequences with length $P$, and the mean energy of each segment can be calculated by
@@
Figure 9: Bouding box of T0011D00S00L0 at SNR = 15 dB. (a) Original Bounding Box. (b) With time domain parameters correction.
Figure 10: Bouding box T0001D00S00L1 at SNR = -9 dB. (a) Original Bounding Box. (b) With frequency domain correction. (c) With time domain correction.
@@
IV Broadcast Frame Decoding Algorithm
For a broadcast frame transmitted at $f_{i}$, the bitstream is obtained after performing time and frequency synchronization, which is then decoded to recover the payload information.
IV-A Time and Frequency Synchronization
  1. 1.
    Frequency Coarse Synchronization: The signal frequency is shifted from $f_{i}$ to near-zero, then $\hat{x}(n)$ will pass through a low-pass filter and be resample to obtain $\dot{x}(m)$.
  2. 2.
    Time Coarse Synchronization: The start of OFDM symbols can be determined by the peak of auto-correlated results of CP, which can be expressed as
    $\gamma(m)=\frac{{{\left|\sum\limits_{n=1}^{{N_{\text{cp}}}}{\dot{x}(m+n){{{\dot{x}}}^{*}}(m+n+N)}\right|}^{2}}}{\left[\sum\limits_{n=1}^{{N_{\text{cp}}}}{\dot{x}(m+n){{{\dot{x}}}^{*}}(m+n)}\right]\left[\sum\limits_{n=1}^{{N_{\text{cp}}}}{\dot{x}(m+n+N){{{\dot{x}}}^{*}}(m+n+N)}\right]}.$ (18)
  3. 3.
    Frequency Fine Synchronization: The frequency offset $\varepsilon$ can be estimated by $-\frac{\theta N}{2\pi h}$, where $\theta$ denote the phase angle at the peak of cross-correlation results based on CP or ZC sequences, and $h$ can be either $N_{\text{cp}}$ or $N$. Then we have
    $\ddot{x}(n)=\dot{x}(n)\text{e}^{-\frac{\text{j}2\pi t}{N}\varepsilon},t=1,2,\ldots,T.$ (19)
  4. 4.
    Time Fine Synchronization: The start of symbols will be re-estimated based on $\ddot{x}(n)$ and (18). Subsequently, CP will be removed, and OFDM symbols are demodulated and quantized to obtain bitstream.
IV-B Bitstream Decode
The bitstream is descrambled based on the Gold sequences and subsequently processed by a Turbo decoder to recover the payload information[14],[15],[52], and the cyclic redundancy check (CRC) value contained in each frame will confirms the correctness of decoded result. In addition, the remaining data can also verify the correctness of decoding results. Since the datasets come from Zhejiang University, the latitude and longitude in the decoding results should match it. The first three symbols of the serial number corresponding to different types of drones are different and fixed, thus the decoding result should also be consistent with the regulations.
V Numerical Results
V-A Simulation Setup